The electric current existing in a discharge tube is 2.0 μA. How much charge is transferred across a cross-section of the tube in 5 minutes?

Grade:10

1 Answers

Kevin Nash
9 years ago
Sol. I’ = 2 μA, t = 5 min = 5 * 60 sec. q = i t = 2 * 10^–6 * 5 * 60 = 10 * 60 * 10^–6 c = 6 * 10^–4 c